# Generating Hypotheses of Dynamic Causal Graphs in Neuroscience: Leveraging Generative Factor Models of Observed Time Series

### v1

@@ A.2 Noisy Classification vs Generative Task Information Requirements
We now provide a theoretical argument as to why it may be necessary to provide additional system history to (and even drastically limit the modeling capabilities of) the REDCLIFF(-S) state model. Our argument centers around the implications of the $\tau_{\mathrm{in}}$ and $\tau_{\mathrm{cl}}$ parameters (Eq. 4), and is thus related to long-standing prior work investigating the effects of similar parameters on state-space models of dynamical systems (Kugiumtzis, 1996). Here, we argue that the task of predicting the state of a signal requires more historical time steps than is required to generate the same time series in the presence of nontrivial noise, at least for some dynamical systems. We now present a summary of our theoretical argument, followed by a proof sketch, and finally by a discussion of the theoretical implications.
@@ A.2.2 Lemma 1: Statement and Proof Sketch
Lemma 1 To learn an accurate, unbiased state model $g_{\theta}$ mapping recorded signals with nontrivial noise to states of $\Phi$ (i.e. $g_{\theta}:\mathbb{R}^{D\times T}\rightarrow\mathbb{N}_{\leq N}$), it must be that $T>\tau$.
Proof Sketch Lemma 1 follows from Bayes’ Theorem, which we use to show that two windows of length $\tau$ sampled from a distribution featuring nontrivial noise can’t be readily distinguished by an unbiased model $g_{\theta}$.
Suppose the recording $\mathbf{x}\in\mathbb{R}^{D\times T}$ with $T<\tau$ is randomly sampled from $\Phi$ according to a uniform distribution over all $N$ states. By way of contradiction, assume we have an unbiased classifier (i.e. the REDCLIFF-S state model $g_{\theta}$) for which $p(f_{i}|\mathbf{X})\neq p(f_{j}|\mathbf{X})$.
Notice that since both $\mathbf{A}_{i}$ and $\mathbf{A}_{j}$ are invertible, we have both that
$\mathbf{X}_{:,T-\tau}=f_{i}^{-1}(\mathbf{X}_{:,T})=\mathbf{A}_{i}^{-1}\mathbf{X}_{:,T}\sim p(\mathbf{X}|f_{i})$
as well as that
$\mathbf{X}_{:,T-\tau}=f_{j}^{-1}(\mathbf{X}_{:,T})=\mathbf{A}_{j}^{-1}\mathbf{X}_{:,T}\sim p(\mathbf{X}|f_{j}).$
@@ A.2.3 Discussion
Lemma 1 has important implications for training REDCLIFF-based models. Effectively, once the state model $g_{\theta}$ in a REDCLIFF model has enough capacity to accurately distinguish between states, Lemma 1 indicates that $g_{\theta}$ may also have access (by necessity) to at least as much information as that required to forecast the evolution of $\Phi$ (assuming the proper noise profile) in one or more system states, at least for restricted (yet simple and arguably common) classes of temporal systems. These findings indicate that it may be necessary at times to restrict the modeling capabilities of $g_{\theta}$ to ensure proper delineation of tasks within a REDCLIFF-S model.
@@
Here we formally define broadcast multiplication - the linear operation used to combine outputs of our generative factors.
@@ Appendix B Methodological Details
As mentioned in the main paper, we implement each factor of the REDCLIFF-S models in this work as a “component-wise Multi-Layer Perceptron” or “cMLP” (Tank et al., 2021); here we briefly summarize the cMLP architecture and refer the reader to the original paper for more details. Essentially, the cMLP architecture is a 1-D convolutional neural network that treats the first layer of weights as an information filter mapping prior variable values of a dynamic system to the next predicted value of a single variable. Since the architecture allows for a set of historical system observations to be input, the first layer of a cMLP comes to represent a set of estimated, time-lagged Granger causal connections. One can view the magnitude of each weight in the first layer of parameters as the strength of a Granger causal connection from one of the input variables to the predicted variable. By instantiating a different cMLP for each variable in a dynamical system, one can learn a sequence of time-lagged, full adjacency matrices mapping prior variable values to future variable values. We found this architecture to be flexible for our purposes, and it proved at least as effective in predictive tasks as other architectures we tested in preliminary experiments (namely the cLSTM (Tank et al., 2021) and DGCNN (Song et al., 2018) architectures).
@@ B.2 Standardizing Causal Graph Estimates
To facilitate fair comparison between both baseline and our REDCLIFF-based algorithms, we needed to standardize the form in which each algorithm’s estimated Granger causal graphs were presented. This was non-trivial, since some of the algorithms incorporated a lag-dimension into their causal graph estimates (esp. those based on the cMLP architecture from Tank et al. (2021)), while others (esp. DYNOTEARS from Pamfil et al. (2020)) treated each feature measured from a channel - be it a historical scalar value or a power-spectral feature - as an individual node in the estimated graph, while yet others (esp. dCSFA-NMF in Gallagher et al. (2021) and REDCLIFF-based models) learned multiple causal graph factors.
The standard we chose to adopt was that each estimated (and true) Granger causal graph would need to be presented as a simple adjacency matrix; i.e. if there were $\mathrm{n_{c}}$ variables in the recorded system, then the true/estimated graph(s) from each system/model were compressed to lie within $\mathbb{R}^{\mathrm{n_{c}}\times\mathrm{n_{c}}}$. In practice, this meant that graphs involving lags and/or node features had their representations compressed by summing over the lag and/or feature values corresponding to each node.
If an algorithm was only capable of estimating a single graph for a system governed by $\mathrm{n_{k}}>1$ graphs, the estimated graph was copied $\mathrm{n_{k}}$ times and compared to each of the $m$ true causal graphs, with results averaged across each comparison.
@@ B.3 REDCLIFF-S Stopping Criteria and Model Selection
Here we describe how we arrived at the stopping criteria for REDCLIFF-S model training and how it was used in model selection. Our stopping criteria (Equation 10) has three main components: a forecasting term, a state classification term, and an estimated Granger causal graph similarity term. Intuitively, the two MSE terms (forecasting and state classification) are in the stopping criteria to ensure our performance on the two main subtasks of our REDCLIFF-S models (predicting factor effect and relevance; see Figure 1A) is being optimized.
The third cosine similarity-based (estimated Granger causal graph similarity) term is included for several reasons. For one thing, the very premise of REDCLIFF-S is that the target system being modeled has different causal relationships over time; hence, our stopping criteria implicitly assumes that component factors ought to be different by minimizing this estimated Granger causal graph similarity term. A potentially useful side-effect of minimizing similarity between factors is that it reduces the size and complexity of our REDCLIFF-S model.
In addition to these intuitive motivations for minimizing cosine similarity between factors, we also found this cosine similarity-based term was correlated with the value of various graph similarity metrics comparing factor estimates to their ground truth representations over the course of training (for example, see Supplemental Table 4). Other interesting phenomena include our finding that troughs in our mean cosine similarity curve coincided with peaks in the DeltaCon0 similarity (a measure of functional similarity between two graphs (Koutra, 2015)) between estimated and true causal graphs as well as the onset of instabilities in the ROC-AUC curves comparing estimates to ground truth. These observations (along with similar ones from many other experiments) demonstrate why this penalty for the sum of cosine similarity between factor estimates served as a reasonable proxy to determine when to stop training.
@@
One open question is how to balance the significant differences in variation between the forecasting and state prediction penalties on one hand and the significantly less-variable cosine similarity term on the other. In selecting the final model for our D4IC experiment, we found that weighting the cosine similarity penalty by $60\rho$ (instead of just $\rho$) in our selection criteria balanced the three variation scales to all lie within roughly the same order of magnitude. More sophisticated balancing schemes would likely improve results.
@@ B.4 Discussion: Limitations and Future Work
Most limitations of REDCLIFF-S arise as trade-offs for scientific utility. For example, the set of REDCLIFF-S factors has many hyperparameters which can be costly to tune; this cost restricted us to tuning on a single repeat of our 12-node, 33-edge, 3-factor system in Section 4.2, and may have limited REDCLIFF-S’ measured performance.
@@ B.5 Comparing State Prediction Error with 1-Factor Models
To test the ability of the REDCLIFF-S algorithm to dynamically adjust factor weightings at each time step in preliminary experiments, we compared the mean-squared error of each state label prediction $\hat{\mathbf{y}}$ made by REDCLIFF-S models against a ‘Naïve 1-Factor Model’ state prediction. For this, we adopted the convention that a state label prediction made by a 1-Factor model (such as the cMLP and cLSTM models by (Tank et al., 2021) or the DGCNN method by (Song et al., 2018)) would be equivalent to predicting that all system factors (represented by the model’s single factor) would always be present; hence, $\hat{\mathbf{y}}=\mathbf{1}$ for any input $\mathbf{x}_{in}$. The difference in MSE between these naïve predictions and the true label $\mathbf{y}$ was taken with the MSE obtained by the REDCLIFF-S models’ predictions across many samples and then averaged.
@@ Appendix C Additional Experimental Results
C.1 Pair-wise Optimal F1-Score Improvement of REDCLIFF-S on Synthetic Systems’ 6-2-2 Data
@@
Figure 6: Pair-wise optimal f1-Score Improvement of REDCLIFF-S on the Synthetic Systems 6-2-2 experiment.
@@ C.2 Additional D4IC Results
We also report ROC-AUC statistics corresponding to D4IC-trained REDCLIFF-S, cMLP-v2, and dCSFA-NMF models in Table 5.
@@
Figure 9: Summary plot of pair-wise optimal f1-score improvement of REDCLIFF-S across Synthetic Systems experiments.
@@ C.3 Additional Synthetic Systems Results
We give additional results related to Table 2 in Supplemental Table 7. Note that between the two tables, the top performing baselines - usually Regime-PCMCI (Saggioro et al., 2020) and SLARAC (Weichwald et al., 2020) - appear to ‘trade-off’ between performance on metrics such as optimal f1-score and structural hamming distance (SHD) versus performance on the AID metrics. In contrast, REDCLIFF-S does not exhibit this trade-off, regularly attaining competitive performance (statistically similar to the best or second-best performances) across all metrics.
@@
With regards to optimal f1-score performance, we provide a summary plot of these results in Supplemental Figure 9 and a visualization comparing true to estimated factors in Supplemental Figure 10. Optimal f1-score and ROC-AUC score results are grouped by systems rated with similar complexity scores. Scores of $\mathfrak{C}\leq 7.0$ received a ‘Low’ categorization; results on systems with this rating can be found in Supplementary Figures 12 through 16. Scores which satisfied $7.0<\mathfrak{C}\leq 13.0$ were categorized as ‘Moderate’ complexity; results on systems with this rating can be found in Supplementary Figures 18 through 22. Scores were categorized as ‘High’ complexity if they satisfied $\mathfrak{C}>13.0$; results on systems with this rating can be found in Supplementary Figures 24 through 24.
@@ C.4 Additional Region-Averaged TST 100 Hz Results
We provide a visualization of our approach to model selection (specifically with regard to determining the number of REDCLIFF-S factors) in Supplemental Figure 26. Plots of the supervised adjacency matrices learned by our REDCLIFF-S models trained on the Region-Averaged TST 100 Hz dataset can be found in Supplemental Figures 26 through 28. We provide additional plots of the differences between the average Open Field (OF) and Tail Suspended (TS) supervised factors in Supplemental Figure 30 and between the average TS and Home Cage (HC) supervised factors in Supplemental Figure 30.
@@ C.5 Region-Averaged Social Preference 100 Hz Results
In addition to the TST dataset, we also applied our REDCLIFF-S method to analyze the Social Preference (SP) dataset described in Mague et al. (2022). It should be noted that our analysis here is preliminary, and was not as thorough as our analysis of the TST data due to time constraints. Supplemental Figure 31 provides a visualization of how we selected the number of factors for the REDCLIFF-S models trained on different folds of the SP data; note that the selection criteria does not decrease monotonically as the number of factors increases (as it did on the TST data), suggesting that a more detailed search of REDCLIFF-S’ hyperparameter settings may yield more accurate results. Plots of the supervised adjacency matrices learned by our REDCLIFF-S models trained on the Region-Averaged SP 100 Hz dataset can be found in Supplemental Figures 33 and 33.
We provide an additional plot of the mean normalized difference between Social Preference (SP) and Object Preference (OP) supervised factors in Supplemental Figure 34. While several of the connections that REDCLIFF-S found do appear to have been published previously (compare with Figure 4A in Mague et al. (2022)), we highlight the fact that the network identified by REDCLIFF-S seems to break from previously published results by including the Hippocampus in multiple network edges. Interestingly, Mague et al. (2022) report that elevated power in certain frequency bands recorded from the Hippocampus was implicated in social behavior, but they did not seem to detect any coherence between activity in the Hippocampus versus other regions. Given that our REDCLIFF-S method is designed to estimate both linear and nonlinear granger causal relationships whereas the dCSFA-based technique employed by Mague et al. (2022) focuses exclusively on linear causal relationships, these findings may suggest that the Hippocampus shares nonlinear causal relationships with other brain regions during social activity. More follow up work would need to be done to confirm this hypothesis, but we are excited to see that REDCLIFF-S does appear to be identifying plausible, novel hypotheses.

### v2

@@ A.2 Noisy Classification vs Generative Task Information Requirements
We now provide a theoretical argument as to why it may be necessary to provide additional system history to the REDCLIFF(-S) state model $g_{\theta}$, and even to drastically limit its modeling capabilities (e.g. via the invertible $g_{\alpha}$ and $g_{y}$ functions of Section 3.5). Our argument centers on the implications of the parameters $\tau_{\mathrm{in}}$ and $\tau_{\mathrm{cl}}$ (Eq. 4), and is thus related to long-standing prior work investigating the effects of similar parameters on state-space models of dynamical systems (Kugiumtzis, 1996). Here, we argue that the task of predicting the state of a signal requires more historical time steps than is required to generate the same time series in the presence of nontrivial noise, at least for some dynamical systems. We now present a summary of our theoretical argument, followed by a proof sketch, and finally a discussion of the theoretical implications.
@@ A.2.2 Lemma 1: Statement and Proof Sketch
Lemma 1 To learn an accurate and unbiased state model $g_{\theta}$ mapping recorded signals with nontrivial noise to states of $\Phi$ (i.e., $g_{\theta}:\mathbb{R}^{D\times T}\rightarrow\mathbb{N}_{\leq N}$), it must be true that $T>\tau$.
Proof Sketch Lemma 1 follows from Bayes’ Theorem, which we use to show that two windows of length $\tau$ sampled from a distribution featuring nontrivial noise cannot be readily distinguished by an unbiased model $g_{\theta}$.
Suppose the recording $\mathbf{x}\in\mathbb{R}^{D\times T}$ with $T<\tau$ is randomly sampled from $\Phi$ according to a uniform distribution over all $N$ states. By way of contradiction, assume we have an unbiased classifier (i.e., the REDCLIFF-S state model $g_{\theta}$) for which $p(f_{i}|\mathbf{X})\neq p(f_{j}|\mathbf{X})$.
Notice that since both $\mathbf{A}_{i}$ and $\mathbf{A}_{j}$ are invertible, we have both that $\mathbf{X}_{:,T-\tau}=f_{i}^{-1}(\mathbf{X}_{:,T})=\mathbf{A}_{i}^{-1}\mathbf{X}_{:,T}\sim p(\mathbf{X}|f_{i})$ as well as that $\mathbf{X}_{:,T-\tau}=f_{j}^{-1}(\mathbf{X}_{:,T})=\mathbf{A}_{j}^{-1}\mathbf{X}_{:,T}\sim p(\mathbf{X}|f_{j}).$
@@ A.2.3 Discussion
Lemma 1 has important implications for training REDCLIFF-based models. Effectively, once the state model $g_{\theta}$ in a REDCLIFF model has enough capacity to accurately distinguish between states, Lemma 1 indicates that $g_{\theta}$ may also have access (by necessity) to at least as much information as that required to forecast the evolution of $\Phi$ (assuming the proper noise profile) in one or more states of the system, at least for restricted (yet simple and arguably common) classes of temporal systems. These findings indicate that it may be necessary to restrict the modeling capabilities of $g_{\theta}$ to ensure the proper delineation of tasks within a REDCLIFF-S model.
@@
Here we formally define broadcast multiplication, the linear operation used to combine outputs of our generative factors.
@@ Appendix B Methodological Details
As mentioned in the main paper, we implement each factor of the REDCLIFF-S models in this work as a “component-wise Multi-Layer Perceptron” or “cMLP” (Tank et al., 2021); here we briefly summarize the cMLP architecture and refer the reader to the original paper for more details. Essentially, the cMLP architecture is a 1D convolutional neural network that treats the first layer of weights as an information filter mapping prior variable values of a dynamic system to the next predicted value of a single variable. Since the architecture allows for a set of historical system observations to be input, the first layer of a cMLP comes to represent a set of estimated, time-lagged Granger causal connections. One can view the magnitude of each weight in the first layer of parameters as the strength of a Granger causal connection from one of the input variables to the predicted variable. By instantiating a different cMLP for each variable in a dynamical system, one can learn a sequence of time-lagged, full adjacency matrices mapping prior variable values to future variable values. We found this architecture to be flexible for our purposes and it proved at least as effective in predictive tasks as other architectures we tested in preliminary experiments (namely the cLSTM (Tank et al., 2021) and DGCNN (Song et al., 2018) architectures).
@@ B.2 Standardizing Causal Graph Estimates
To facilitate fair comparison between both baseline and our REDCLIFF-based algorithms, we needed to standardize the form in which each algorithm’s estimated Granger causal graphs were presented. This was nontrivial, since some of the algorithms incorporated a lag-dimension into their causal graph estimates (esp. those based on the cMLP architecture from Tank et al. (2021)), while others (esp. DYNOTEARS from Pamfil et al. (2020)) treated each feature measured from a channel - be it a historical scalar value or a power spectral feature - as an individual node in the estimated graph, while yet others (esp. dCSFA-NMF in Gallagher et al. (2021) and REDCLIFF-based models) learned multiple causal graph factors.
The standard we chose to adopt was that each estimated (and true) Granger causal graph would need to be presented as a simple adjacency matrix; that is, if there were $\mathrm{n_{c}}$ variables in the recorded system, then the true/estimated graph(s) from each system/model were compressed to lie within $\mathbb{R}^{\mathrm{n_{c}}\times\mathrm{n_{c}}}$. In practice, this meant that graphs involving lags and/or node features had their representations compressed by summing over the lag and/or feature values corresponding to each node.
If an algorithm could only estimate a single graph for a system governed by $\mathrm{n_{k}}>1$ graphs, the estimated graph was copied $\mathrm{n_{k}}$ times and compared to each of the $m$ true causal graphs, with results averaged across each comparison.
@@ B.3 REDCLIFF-S Stopping Criteria and Model Selection
Here we describe how we arrived at the stopping criteria for REDCLIFF-S model training and how it was used in model selection. Our stopping criteria are summarized in Equation 10, which has three main components: a forecasting term, a state classification term, and an estimated Granger causal graph similarity term. Intuitively, the two MSE terms (forecasting and state classification) are in the stopping criteria to ensure that our performance on the two main subtasks of our REDCLIFF-S models (predicting factor effect and relevance; see Figure 1A) is being optimized.
The third cosine similarity-based term (estimated Granger causal graph similarity) is included for several reasons. For one thing, the very premise of REDCLIFF-S is that the target system being modeled has different causal relationships over time; hence, our third term implicitly assumes that factors should be different by minimizing this estimated Granger causal graph similarity term. A potentially useful side effect of minimizing similarity between factors is that it reduces the size and complexity of our REDCLIFF-S model.
In addition to these intuitive motivations for minimizing cosine similarity between factors, we also found this cosine similarity-based term was correlated with the value of various graph similarity metrics comparing factor estimates with their ground truth representations over the course of training (for example, see Supplemental Table 4). Other interesting phenomena include our finding that troughs in our mean cosine similarity curve coincided with peaks in the DeltaCon0 similarity (a measure of functional similarity between two graphs (Koutra, 2015)) between estimated and true causal graphs, as well as the onset of instabilities in the ROC-AUC curves when comparing estimates to ground truth. These observations (along with similar ones from many other experiments) demonstrate why this penalty for the sum of cosine similarity between factor estimates served as a reasonable proxy to determine when to stop training.
@@
One open question is how to balance the significant differences in variation between the forecasting and state prediction penalties on the one hand and the significantly less-variable cosine similarity term on the other. In selecting the final model for our D4IC experiment, we found that weighting the cosine similarity penalty by $60\rho$ (instead of just $\rho$) in our selection criteria balanced the three variation scales to all roughly lie within the same order of magnitude. More sophisticated balancing schemes would likely improve results.
@@ B.4 Discussion: Limitations and Future Work
Most limitations of REDCLIFF-S arise as trade-offs for scientific utility. For example, the set of REDCLIFF-S factors has many hyperparameters that can be costly to tune; this cost restricted us to tuning on a single repeat of our 12-node, 33-edge, 3-factor system in Section 4.2, and may have limited REDCLIFF-S’ measured performance.
@@ B.5 Comparing State Prediction Error with 1-Factor Models
To test the ability of the REDCLIFF-S algorithm to dynamically adjust factor weightings at each time step in preliminary experiments, we compared the mean squared error of each state label prediction $\hat{\mathbf{y}}$ made by REDCLIFF-S models against a ‘Naïve 1-Factor Model’ state prediction. For this, we adopted the convention that a state label prediction made by a 1-factor model (such as the cMLP and cLSTM models by (Tank et al., 2021) or the DGCNN method by (Song et al., 2018)) would be equivalent to predicting that all system factors (represented by the model’s single factor) would always be present; hence, $\hat{\mathbf{y}}=\mathbf{1}$ for any input $\mathbf{x}_{in}$. The difference in MSE between these naïve predictions and the true label $\mathbf{y}$ was taken with the MSE obtained by the predictions of the REDCLIFF-S models across many samples and then averaged.
@@ Appendix C Additional Experimental Results
C.1 Pairwise Optimal F1-Score Improvement of REDCLIFF-S on Synthetic Systems’ 6-2-2 Data
@@
Figure 6: Pairwise optimal f1-Score Improvement of REDCLIFF-S on the Synthetic Systems 6-2-2 experiment.
@@ C.2 Additional D4IC Results
We also report ROC-AUC statistics corresponding to REDCLIFF-S, cMLP-v2, and dCSFA-NMF models trained on the D4IC dataset in Table 5.
@@
Figure 9: Summary plot of pairwise optimal f1-score improvement of REDCLIFF-S across Synthetic Systems experiments.
@@ C.3 Additional Synthetic Systems Results
We give additional results related to Table 2 in Supplemental Table 7. Note that between the two tables, the top performing baselines - usually Regime-PCMCI (Saggioro et al., 2020) and SLARAC (Weichwald et al., 2020) - appear to ‘trade off’ between performance on metrics such as optimal f1-score and structural hamming distance (SHD) versus performance on the AID metrics. In contrast, REDCLIFF-S does not exhibit this trade-off, regularly attaining competitive performance (statistically similar to the best or second-best performance) across all metrics.
@@
Regarding optimal f1-score performance, we provide a summary plot of these results in Supplemental Figure 9 and a visualization comparing true to estimated factors in Supplemental Figure 10. Optimal f1-score and ROC-AUC score results are grouped by systems rated with similar complexity scores. Scores of $\mathfrak{C}\leq 7.0$ received a ‘Low’ categorization; results on systems with this rating can be found in Supplementary Figures 12 through 16. Scores that satisfied $7.0<\mathfrak{C}\leq 13.0$ were categorized as ‘Moderate’ complexity; results on systems with this rating can be found in Supplementary Figures 18 through 22. Scores were categorized as ‘High’ complexity if they satisfied $\mathfrak{C}>13.0$; the results on systems with this rating can be found in Supplementary Figures 24 through 24.
@@ C.4 Additional Region-Averaged TST 100 Hz Results
We provide a visualization of our approach to model selection (specifically with respect to determining the number of REDCLIFF-S factors) in Supplemental Figure 26. Plots of the supervised adjacency matrices learned by our REDCLIFF-S models trained on the Region-Averaged TST 100 Hz dataset can be found in Supplemental Figures 26 through 28. We provide additional plots of the differences between the average Open Field (OF) and Tail Suspended (TS) supervised factors in Supplemental Figure 30 and between the average TS and Home Cage (HC) supervised factors in Supplemental Figure 30.
@@ C.5 Region-Averaged Social Preference 100 Hz Results
In addition to the TST dataset, we also applied our REDCLIFF-S method to analyze the Social Preference (SP) dataset described in Mague et al. (2022). It should be noted that our analysis here is preliminary and was not as thorough as our analysis of the TST data due to time constraints. Supplemental Figure 31 provides a visualization of how we selected the number of factors for the REDCLIFF-S models trained on different folds of the SP data; note that the selection criteria history does not decrease monotonically as the number of factors increases (as it did on the TST data), suggesting that a more detailed search of REDCLIFF-S’ hyperparameter settings may yield more accurate results. Plots of the supervised adjacency matrices learned by our REDCLIFF-S models trained on the Region-Averaged SP 100 Hz dataset can be found in Supplemental Figures 33 and 33.
We provide an additional plot of the mean normalized difference between Social Preference (SP) and Object Preference (OP) supervised factors in Supplemental Figure 34. While several of the connections found by REDCLIFF-S do appear to have been published previously (compare with Figure 4A in Mague et al. (2022)), we highlight the fact that the network identified by REDCLIFF-S seems to break from previously published results by including the Hippocampus in multiple network edges. Interestingly, Mague et al. (2022) report that elevated power in certain frequency bands recorded in the Hippocampus was implicated in social behavior, but they did not seem to detect any coherence between activity in the Hippocampus and other regions. Given that our REDCLIFF-S method is designed to estimate both linear and nonlinear granger causal relationships, whereas the dCSFA-based technique employed by Mague et al. (2022) focuses exclusively on linear causal relationships, these findings may suggest that the Hippocampus shares nonlinear causal relationships with other regions of the brain during social activity. More follow-up work would need to be done to confirm this hypothesis, but we are excited to see that REDCLIFF-S appears to be identifying plausible, novel hypotheses.